# The Early Bird Catches the Leak: Unveiling Timing Side Channels in

### v3

@@ 4.1 Analysis on PSA
4.2 Analysis on PNA
Attack overview. In the PNA attack, we note that not all user queries contain sensitive data. An attacker is unlikely to pursue generic requests like “What’s the weather today?”. Instead, they are expected to focus on specific requests more likely to reveal private information. For example, a request such as “Draft a travel plan for my husband and me to Rome in September for a 3-day stay at the Hilton hotel” could expose personal and location details. By identifying such high-risk queries, the attacker can exploit timing side channels to recover private information from other users’ requests. For this purpose, the attacker could compile a list of privacy-related prompts from online sources (Table 2). The goal is to discover the connections with private attributes, e.g., whether a user is linked to a particular medical condition.
@@ 4.2 Analysis on PNA
In practice, users might express the same idea in varied ways and embed diverse private attributes (e.g., personal names, medical conditions) in their queries. Our objective is to determine how these different attributes affect semantic similarity and whether the resulting variations exceed a threshold described in Section 3.3. If so, such differences can be detected, enabling the attacker to determine whether the victim’s request carries particular attributes, even when the victim rephrases the content.
@@
Table 2: Examples of user prompts that contain private attributes.
@@
We tested similarity thresholds from 0.6 to 0.9 (GPTCache’s default is 0.8). 8(a) shows the ROC curves for the positive and negative similarity distributions, revealing a clear separation between the two. At the default threshold of 0.8, for instance, a TPR of 0.95 can be achieved with an FPR below 0.1. We also examined cases where only one private attribute (either $Name_{0}$ or $Medcond_{0}$) matched. Here, the negative group consists of sentences with only one correct private attribute, while the positive group remains the same. 8(b) shows that the semantic distinctions remain substantial: at the default threshold of 0.8, a TPR of 0.85 corresponds to an FPR under 0.1.
(a) The “name” and “medical condition” in the negative group are both different from the positive group.
(b) Either the “name” or “medical condition” in the negative group is different from the positive group.
@@
End-to-end attacks. We consider a typical open-world scenario in which a victim user requests healthcare assistance from an LLM. For instance, the user might submit a query with semantics similar to the template “compose a meeting agenda…” shown in Table 2, but with various names and medical conditions. The user may also send queries unrelated to the targeted request. To simulate this, we model the user’s queries as follows:
@@
Representative requests are those that most closely approximate the rest of the request space. To identify them, we use the distilbert-base-uncased model to generate embeddings and then compute the L2 distance between these embeddings. We then sort the requests by their L2 distances; those with the smallest distances are deemed the most representative and selected as attack requests to maximize coverage. To further expand coverage, we incorporate orthogonal requests—requests that are semantically distinct from one another. This reduces the chance that semantic overlaps among the attacker’s own requests degrade accuracy in identifying victim requests. We classify a cache access as a hit if at least one of the attacker’s requests triggers a cache hit in the timing channel. Although this strategy boosts coverage, it also raises the false positive rate (FPR), necessitating a careful balance.
@@
In the evaluation, each of the 4 victim request types was tested 500 times. In each iteration, a random pair of private attributes in the Type-1 request was selected. We set $\sigma=0.06$ and identified 5 orthogonal attack requests using the proposed greedy search strategy. Section 4.2 summarizes the TPR for true samples, and the separate FPRs for each false sample type when increasing the number of attack requests from 1 to 5. Specifically, we successfully recovered 407 victim requests out of the 500 true samples with a single attack request, achieving a recovery accuracy of 81.4% with an average FPR of 0.045. With 5 attack requests, 477 victim requests are recovered, demonstrating a recovery accuracy of 95.4% with an average FPR of 0.056. We provide a demo for this attack in our website [37].
Table 3: Attack accuracy for the 4 types of victim requests with different number of attack trails.
@@
(TPR) Type 2
(FPR) Type 3
(FPR) Type 3
@@
4.3 Inferring Documents on Commodity LLM
@@
The KV cache can be shared among the same user, within an organization, or even across organizations [48], creating the potential for cross-user information leaks. In our research, we discovered that such leaks are feasible even in remote attack scenarios, particularly when the target LLM processes documents. To demonstrate this, we utilized a document summarization application powered by a commodity LLM API service, where all user requests are processed using the same developer’s API key. We provide an end-to-end demonstration showing how an adversary could infer processed documents from the application through the KV cache side channels. Importantly, an application is not required for cross-user attacks, as the KV cache can be shared across organizations [48]. However, our use of the application highlights the privacy risks inherent to LLM-powered applications, even when they fully comply with the guidelines provided by the LLM service.
@@
The application. We implemented the document summarization application with direct summarization [31] (also known as the stuff approach [4]), using the public Deepseek-chat model as the backend LLM API server. The application was built in accordance with Deepseek’s guideline and operates by first extracting text from user-uploaded PDF files using the pdfplumber package. This text is then formatted into a request and sent to the LLM for summarization. In this setup, documents uploaded by users are included in messages under the “user” role and sent to the Deepseek model, which returns summaries of the documents. Notably, all user inputs are processed under a single Deepseek account, which is typical in such applications. However, this poses a privacy concern because Deepseek’s prompt caching [27] can inadvertently allow cached content to be reused across different users. As a result, a malicious user could potentially infer which documents other users have processed by exploiting timing side channels.
Characterizing the leakage. We experimented with 200 documents of various lengths (approximately 12,000, 18,000, and 24,000 tokens), each saved in the PDF format and derived from segments of the zero_scrolls dataset [70]. Our results revealed distinct latencies between cache hits (where documents had been cached) and cache misses (where documents had not been cached), as shown in Figure 10. In particular, responses to cache hits remained consistently fast across different document lengths, while cache misses grew noticeably slower with larger documents.
@@
Specifically, the victim first submits 100 documents from the interested set. The attacker then probes each of the 200 interested documents once, recording response latencies. Based on a predefined threshold (2.0 seconds in our experiments), the TPR is computed as the fraction of probed documents correctly identified as cache hits (i.e., with latencies below the threshold). Next, the victim uploads 100 additional documents outside the interested set, and the attacker probes the entire set of 200 documents again. The FPR is then calculated as the fraction of documents incorrectly labeled as hits. Our tests produced a TPR of 0.89 and an FPR of 0.05. A demo for the attack is presented in our website [37].
@@
4.4 Measurement Study on Commodity LLMs
@@
Table 4: Summary of KV cache sharing in real world LLM serving systems (date: 08/29/2024).
@@
    We observed a timing difference on 08/29/2024 and reported it to OpenAI. By late December 2024, the timing difference was no longer stable, despite the API indicating that the prompt cache was effective. This may be due to timing obfuscation measures implemented by OpenAI.
Specifically, we first measured the response latencies by sending initial requests that were likely to miss the cache. Then, we sent identical requests multiple times and measured the average latencies for these subsequent requests. To maximize the likelihood of co-locating on the same physical machine and ensuring the requests were cached, we conducted continuous tests within the same time period. If we observed lower latencies in the later requests, this indicated the use of caching mechanisms in the LLM services. With KV cache sharing, the computation of matched prefix tokens during the prefill phase can be ignored. However the output generated during the decoding phase still requires computation and inference, which are influenced by parameters such as temperature, introducing randomness. To verify that the latency reduction was due to KV cache sharing, we also set a high temperature ($0.9$) in the request. We verified whether the LLM produced different responses for each request, with TTFT reductions consistently observed. If it did, this strongly indicated that KV cache sharing was supported, enabling a reduction in TTFT while still allowing for diverse outputs. To minimize the impact of network latency, we sent requests of varying lengths, ranging from 200 to 2,000 tokens. The time difference between cached and uncached responses typically spanned several hundred milliseconds, making it easy to distinguish between the two. Additionally, we observed when the cache is hit, the TTFT remains consistent, regardless of the request length, whereas when the cache is missed, TTFT increases almost linearly as the length of the request grows. As summarized in Table 4, most popular LLM service providers support KV cache sharing in specific scenarios.
Table 5: Native support of semantic caching of popular AI service providers (date: 08/29/2024).
@@
Semantic cache sharing. We manually reviewed the documentation of public cloud AI service providers to verify whether they support semantic cache APIs. As shown in Table 5, semantic caching is supported by major AI platform-as-a-service providers. Notably, even on platforms that do not offer native semantic caching, users can still implement their own solutions or leverage open-source alternatives, such as GPTCache.
5 Mitigations
5.1 Mitigating KV Cache Leakages
@@
Evaluation. To evaluate the effectiveness of the mitigation and reduce the cost of querying the LLM, we conducted a simulation-based experiment. First, we built the classifiers that detect the hits and misses of $K$ tokens for each value of $K$ ($K=1,2,3,4$), following the method outlined in Section 4.1. Since the timing differences become more pronounced as $K$ increases, we reduced the number of samples (i.e., $n$) in multi-sampling, and obtained the corresponding TPRs and FPRs for each classifier. Then we used an oracle to simulate the classifiers by randomly sampling a number and determining whether it fell within the classification range. In this evaluation, we used the same repetitive trails method, dataset and fine-tuned model as described in Section 4.1. The next-token predictor was modified to predict the next $K$ tokens. Section 5.1 presents the token recovery rate and the average number of queries needed to recover 1 token for $K=1,2,3$ and $4$. The results show that as $K$ increases, the attack still achieves a notable recovery rate. The average number of queries for recovered tokens decreases with a larger $K$, as the predictor performs well on easy prompts. However, the overall recovery rate declines because the predictor has more difficulty with harder prompts for a larger $K$. When considering the tokens not recovered after the maximum of 80 allowed guesses, the average number of queries increases.
Table 6: Token recovery results under different numbers of minimum shared tokens.
@@
recovered token #queries
@@
5.2 Mitigating Semantic Cache Leakages
Design. As investigated in Section 4.2, private attributes have a significant impact on the semantic similarity between requests. As a result, the PNA infers private attributes by probing whether a semantically similar request is cached. To mitigate this leakage, we propose a strategy that involves identifying and anonymizing the private attributes present in the requests. This approach not only prevents the leakage of private attributes but also increases the potential for sharing requests across users.
@@ 6 Discussions
7 Related Works
@@
Side channel attacks on LLM. Debenedetti et al. propose system-level side channels within the deep learning lifecycle, such as training data filtering, input preprocessing, output monitoring, and query filtering. These side channels can potentially be exploited to infer the training data or the requests [40]. LLM keystroking attacks [79] are a type of packet analysis based side channels. These attacks exploit the length of response tokens, assuming the attacker has access to encrypted network packets. By comparison, we are the first to study the timing leaks introduced by LLM optimizations, rather than relying on output data or token packet sizes to recover requests. Inspired by our work, a follow-up study by Gu et al. reports a comprehensive measurement analysis of prompt caching in real-world LLMs [48], detecting prompt caching in 8 out of 17 LLM service providers. More recently, Zheng et al. conducted a similar study on LLM timing side channels [94], which however does not feature our optimized search strategy for efficient request recovery.

### v4

@@ IV-A Analysis on PSA
IV-B Analysis on PNA
Attack setup. In the PNA attack, we note that not all user queries contain sensitive data. An attacker is unlikely to pursue generic requests like “What’s the weather today?”. Instead, they are expected to focus on specific requests more likely to reveal private information. For example, a request such as “Draft a travel plan for my husband and me to Rome in September for a 3-day stay at the Hilton hotel” could expose personal and location details. By identifying such high-risk queries, the attacker can exploit timing side channels to recover private information from other users’ requests. For this purpose, the attacker could compile a list of privacy-related prompts from online sources (Table III). The attacker has obtained a list of private attributes, and its goal is to discover connections between entities and these private attributes, e.g., whether a specific user is linked to a particular medical condition.
@@ IV-B Analysis on PNA
In practice, users may express the same intent through varied phrasing and may embed diverse private attributes—such as personal names or medical conditions—within their queries. Our objective is to determine how these different attributes affect semantic similarity and whether the resulting variations exceed a threshold described in Section III-C. If so, such differences can be detected, enabling the attacker to determine whether the victim’s request carries particular attributes, even when the victim rephrases the content.
@@
TABLE III: Examples of user prompts that contain private attributes.
@@
We tested similarity thresholds from 0.6 to 0.9 (the default value in GPTCache is 0.8). Figure9a shows the receiver operating characteristic (ROC) curves for the positive and negative similarity distributions, revealing a clear separation between the two. At the default threshold of 0.8, a TPR of 0.95 can be achieved with an FPR below 0.1. We also examined cases where only one private attribute (either $Name_{0}$ or $Medcond_{0}$) matched. Here, the negative group consists of sentences with only one correct private attribute, while the positive group remains the same. Figure9b shows the semantic distinctions remain substantial: at the default threshold of 0.8, a TPR of 0.85 corresponds to an FPR under 0.1.
(a) (a). The “name” and “medical condition” in the negative group are both different from the positive group.
(b) (b). Either the “name” or “medical condition” in the negative group is different from the positive group.
@@
End-to-end attacks. We consider a typical open-world scenario in which a victim user requests healthcare assistance from an LLM. For instance, the user might submit a query with semantics similar to the template “compose a meeting agenda…” shown in Table III, but with various names and medical conditions. The user may also send queries unrelated to the targeted request. To simulate this, we model the user’s queries as follows:
@@
Representative requests are those that most closely approximate the rest of the request space. To identify them, we use the distilbert-base-uncased [67] model by default to generate embeddings and then compute the L2 distance between these embeddings. We then sort the requests by their L2 distances; those with the smallest distances are deemed the most representative and selected as attack requests to maximize coverage. To further expand coverage, we incorporate orthogonal requests—requests that are semantically distinct from one another. This reduces the chance that semantic overlaps among the attacker’s own requests degrade accuracy in identifying victim requests. We classify a cache access as a hit if at least one of the attacker’s requests triggers a cache hit in the timing channel. Although this strategy boosts coverage, it also raises the FPR, necessitating a careful balance.
@@
In the evaluation, each of the 4 victim request types was tested 500 times. In each iteration, a random pair of private attributes in the Type-1 request was selected. We set $\sigma=0.06$ and identified 5 orthogonal attack requests using the proposed greedy search strategy. Table IV summarizes the TPR for true samples, and the separate FPRs for each false sample type when increasing the number of attack requests from 1 to 5. Specifically, we successfully recovered 407 victim requests out of the 500 true samples with a single attack request, achieving a recovery accuracy of 81.4% with an average FPR of 0.045. With 5 attack requests, 477 victim requests are recovered, demonstrating a recovery accuracy of 95.4% with an average FPR of 0.056. We provide a demo for this attack in our website [62].
TABLE IV: Attack accuracy for the 4 types of victim requests with different number of attack trails.
@@
Type 2
Type 3
Type 3
@@
IV-C Inferring Documents on Commodity LLM
@@
The KV cache can be shared among the same user, within an organization, or even across organizations [31], creating the potential for cross-user information leaks. In our research, we discovered that such leaks are feasible even in remote attack scenarios, particularly when the target LLM processes documents. To demonstrate this, we utilized a document summarization application powered by a commodity LLM API service, where all user requests are processed using the same developer’s API key, with both the victim and the attacker are standard users of the application. We provide an end-to-end demonstration showing how an adversary could infer processed documents from the application through the KV cache side channels. Importantly, such cross-user attacks fundamentally do not require application-specific vulnerabilities, as the KV cache can be shared across organizations [31]. However, our use of the application highlights the privacy risks inherent to LLM-powered applications, even when they fully comply with the guidelines provided by the LLM service.
@@
The victim application. We implemented the document summarization application with direct summarization [68] (also known as the stuff approach [69]), using the public Deepseek-chat model as the backend LLM API server. The application was built in accordance with Deepseek’s guideline and operates by first extracting text from user-uploaded PDF files using the pdfplumber package. This text is then formatted into a request and sent to the LLM for summarization. In this setup, documents uploaded by users are included in messages under the “user” role and sent to the Deepseek model, which returns summaries of the documents. Notably, all user inputs are processed under a single Deepseek account, which is typical in such applications. However, this poses a privacy concern because Deepseek’s prompt caching [70] can inadvertently allow cached content to be reused across different users. As a result, a malicious user could potentially infer which documents other users have processed by exploiting timing side channels.
Characterizing the leakage. We conducted experiments with 200 documents of various lengths (approximately 12,000, 18,000, and 24,000 tokens), each saved in PDF format and derived from segments of the zero_scrolls dataset [71]. Our results revealed distinct latencies between cache hits (where documents had been cached) and cache misses (where documents had not been cached), as shown in Figure 11. In particular, responses to cache hits remained consistently fast across different document lengths, while cache misses grew noticeably slower with larger documents.
@@
Specifically, the victim first submits 100 documents from the interested set. The attacker then probes each of the 200 interested documents once, recording response latencies. Based on a predefined threshold (2.0 seconds in our experiments), the TPR is computed as the fraction of probed documents correctly identified as cache hits (i.e., with latencies below the threshold). Next, the victim uploads 100 additional documents outside the interested set, and the attacker probes the entire set of 200 documents again. The FPR is then calculated as the fraction of documents incorrectly labeled as hits. Our evaluation shows that this attack achieves an average accuracy of 89%, with an average FPR of 0.05. A demo for the attack is presented in our website [62].
@@
IV-D Measurement Study on Commodity LLMs
Specifically, we first measured the response latencies by sending initial requests that were likely to miss the cache. Then, we sent identical requests multiple times and measured the average latencies for these subsequent requests. To maximize the likelihood of co-locating on the same physical machine and ensuring the requests were cached, we conducted continuous tests within the same time period. If we observed lower latencies in the later requests, this indicated the use of caching mechanisms in the LLM services. With KV cache sharing, the computation of matched prefix tokens during the prefill phase can be ignored. However, the output generated during the decoding phase still requires computation and inference, which are influenced by parameters such as temperature, introducing randomness. To verify that the latency reduction was due to KV cache sharing, we deliberately set a high temperature ($0.9$) in the request. This configuration was critical because semantic caching mechanisms typically return identical cached outputs for semantically similar inputs. By introducing substantial randomness in token selection through high temperature, we ensured the LLM would generate diverse responses despite input similarities. We verified whether the LLM produced different responses for each request, with TTFT reductions consistently observed. If it did, this strongly indicated that KV cache sharing was supported, enabling a reduction in TTFT while still allowing for diverse outputs. To minimize the impact of network latency, we sent requests of varying lengths, ranging from 200 to 2,000 tokens. The time difference between cached and uncached responses typically spanned several hundred milliseconds, making it easy to distinguish between the two. Additionally, we observed when the cache is hit, the TTFT remains consistent, regardless of the request length, whereas when the cache is missed, TTFT increases almost linearly as the length of the request grows. As summarized in Table V, most popular LLM service providers support KV cache sharing in specific scenarios.
@@ IV-D Measurement Study on Commodity LLMs
TABLE V: Summary of KV cache sharing in real world LLM serving systems (date: 08/29/2024).
@@
    We observed a timing difference on 08/29/2024 and reported it to OpenAI. By late December 2024, the timing difference was no longer stable, despite the API indicating that the prompt cache was effective.
TABLE VI: Native support of semantic caching of popular AI service providers (date: 08/29/2024).
@@
Semantic cache sharing. We manually reviewed the documentation of public cloud AI service providers to verify whether they support semantic cache APIs. As shown in Table VI, semantic caching is supported by major AI platform-as-a-service providers. Notably, even on platforms that do not offer native semantic caching, users can still implement their own solutions or leverage open-source alternatives, such as GPTCache.
V Mitigations
V-A Mitigating KV Cache Leakages
@@
Evaluation. To evaluate the effectiveness of the mitigation and reduce the cost of querying the LLM, we conducted a simulation-based experiment. First, we built the classifiers that detect the hits and misses of $K$ tokens for each value of $K$ ($K=1,2,3,4$), following the method outlined in Section IV-A. Since the timing differences become more pronounced as $K$ increases, we reduced the number of samples (i.e., $n$) in multi-sampling, and obtained the corresponding TPRs and FPRs for each classifier. Then we used an oracle to simulate the classifiers by randomly sampling a number and determining whether it fell within the classification range. In this evaluation, we used the same repetitive trails method, dataset and fine-tuned model as described in Section IV-A. The next-token predictor was modified to predict the next $K$ tokens. Table VII presents the token recovery rate and the average number of queries needed to recover 1 token for $K=1,2,3$ and $4$. The results show that as $K$ increases, the attack still achieves a notable recovery rate. For successfully recovered tokens, the average number of queries decreases with larger $K$ values, as the predictor quickly identifies tokens in easier-to-predict prompts. However, the overall recovery rate declines because the predictor increasingly struggles with harder-to-predict tokens at higher K values. When accounting for tokens not recovered within the maximum 80 allowed guesses, the average query cost per token recovery increases significantly.
TABLE VII: Token recovery results under different numbers of minimum shared tokens.
@@
recovered token
#queries
@@
V-B Mitigating Semantic Cache Leakages
Design. As investigated in Section IV-B, private attributes have a significant impact on the semantic similarity between requests. As a result, the PNA infers private attributes by probing whether a semantically similar request is cached. To mitigate this leakage, we propose a strategy that involves identifying and anonymizing the private attributes present in the requests. This approach not only prevents the leakage of private attributes but also increases the potential for sharing requests across users.
@@ VI Discussions
Real-world attacks. In real world attacks, especially targeting online LLM services, the network latency will bring additional noises to the timing. The noise effect could be reduced with more attack trials. Besides, it has been shown possible to mount cache attacks from the remote network [41, 42].
VII Related Works
@@
Side channel attacks on LLM. Debenedetti et al. propose system-level side channels within the deep learning lifecycle, such as training data filtering, input preprocessing, output monitoring, and query filtering. These side channels can potentially be exploited to infer the training data or the requests [94]. LLM keystroking attacks [95] are a type of packet analysis based side channels. These attacks exploit the length of response tokens, assuming the attacker has access to encrypted network packets. Besides, Carlini et al. investigated the privacy risks of timing attacks in multi-turn interactions with LLM chatbots [96]. By comparison, we are the first to study the timing leaks introduced by LLM serving system optimizations, rather than relying on output data or token packet sizes to recover requests.